OpenStudy (anonymous):

tickets sales have gone up 10% reaching 880 tickets. how many ticket were sold in the last season.

OpenStudy (anonymous):

Let x be the number of tickets in the last season. x+10%x=880 x+0.1x=880 1.1x=880 x=800 Therefore, there were 800 tickets sold in the last season.

OpenStudy (anonymous):

i would like to leave x out of it

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

putting the x in just messes me up u loss me with 1.1x=880 x=800 shouldn't it just be multipy and divide

OpenStudy (anonymous):

okay, just divide 880 by 1.1 to get the answer.

OpenStudy (anonymous):

why is it 1.1

OpenStudy (anonymous):

why do we add 1....this is not a joke I really don't understand

OpenStudy (anonymous):

1.1 is from adding 1 to 10%(10%is equal to 0.1) the answer is 1.1.

OpenStudy (anonymous):

okay, please go back to the equation that i gave you. it will tell you why we have 1.

OpenStudy (anonymous):

have you done pre algebra?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ah, that is why, actually, we are translating word problems to mathematical equations here.

OpenStudy (anonymous):

that is the problem for me....just don't understand....I can do percentage but these are kicking my butt

OpenStudy (anonymous):

okay let us do the percentage then

OpenStudy (anonymous):

P=br

OpenStudy (anonymous):

what is br

OpenStudy (anonymous):

so, try to analyze the problem. It will give us the idea that the tickets sold is 100%. If we are going to have the equation for the tickets sold for the present, it is number of tickets sold + 10% of it, right? so, we can have P=880 r=100%+10%, or r=110% P=br 880=b(110%) so, we are going to divide 880 by 110%(or this is 1.1 in decimal form) As we divide, we will get 800, which is the number of tickets sold in the last season.

OpenStudy (anonymous):

wow I would of never got that... thanks..you must be very good I see you answering a lot of questions.

OpenStudy (anonymous):

thank you and you are welcome. kindly click the good answer for me.lol..hehehe